# The concept of minimal dissipation and the identification of work in autonomous systems: A view from classical statistical physics

### v1

@@ VI Projection operator for quantum systems
It is not straightforward to transfer the reasoning for classical systems to quantum systems in a universally valid manner. The meaning of nonlinear combinations of the relevant observables depends on their order. Relevant observables are represented by operators which do not commute in general. For a single relevant observable $A$ we could use its spectral decomposition to define the projector. Suppose $A$ has a discrete spectrum
@@
Due to orthogonality we have $\Pi_{j}\Pi_{k}=\delta_{jk}\Pi_{j}$, much the same as in eq. (25). Both $\mathcal{P}$ and $\mathcal{P}^{\dagger}$ are idempotent. The projected $\mathcal{P}^{\dagger}\rho$ is still a density matrix. It can be interpreted as the projection on the equilibrium state after a non-selective measurement of $A$. By construction, the expectation value of $A$ is preserved, i.e., $\text{Tr}(A\rho)=\text{Tr}(A\mathcal{P}^{\dagger}\rho)$. At first glance this seems like a promising pathway, but there are several pitfalls.
@@
Further, this construction does not yield an equation of motion for the reduced density matrix $\rho_{S}$ of a composite system. For a general density matrix $\rho_{S}=\text{Tr}_{E}(\rho)\neq\text{Tr}_{E}(\mathcal{P}^{\dagger}\rho)$. Consider the situation where the relevant observable is a tensor product of a system operator and the environmental identity, i.e. $A=A_{S}\otimes\mathbb{I}_{E}$ where $A_{S}$ is an operator which acts in $\mathcal{H}_{S}$. With the spectral decomposition of $A_{S}=\sum_{j}a_{j}\Pi_{j}$ and $\sum_{j}\Pi_{j}=\mathbb{I}_{S}$, we can just obtain an EoM for the part of $\rho_{S}$ which commutes with $A_{S}$.
In addition, we could also define (52) with a different operator order in the numerator and a different distribution of the weight $\varrho$. The latter is due to the fact that there is no unique quantum analogue to the classical scalar product (26). Instead of the straightforward choice $\text{Tr}(\varrho_{\beta}X^{\dagger}Y)$ we could switch the order or use the symmetrized variant $\frac{1}{2}\text{Tr}(\varrho_{\beta}(X^{\dagger}Y+Y^{\dagger}X))$ as used in ref. [38]. There is an entire class of scalar products
@@
giving a weight to the commutativity property with $H$ [39]. A common choice in quantum statistical mechanics is to average over $\alpha$
@@
This scalar product is known under various names like Mori scalar product, Kubo’s canonical correlation, Bogoliubov inner product or Duhamel two point function [26, 27, 40].
@@
The advantage of this form is, that
@@
where $\Sigma_{S}$ maps system operators $X_{S}\in\mathcal{H}_{S}$ to system operators: $\Sigma_{S}X_{S}=\text{Tr}_{E}(\Sigma X_{S})$. With these definitions we formally have the desired $\mathcal{P}X_{S}=X_{S}$ and $\text{Tr}_{E}(\mathcal{P}^{\dagger}\rho)=\text{Tr}_{E}(\rho)=\rho_{S}$, i.e. all observables of the system are relevant and the relevant density yields the reduced system state if we trace over the environmental degrees of freedom.
@@
If the action of $\Sigma_{S}$ can be represented analogous to $\Sigma$ with $\varrho_{\beta}$ replaced by $\varrho_{S,\beta}$ in eq. (56), a relation similar to eq. (57) holds such that $-\Sigma_{S}[X_{S},H_{S}^{\ast}]=\frac{1}{\beta}[X_{S},\rho_{S,\beta}]$. Accordingly we would obtain the desired commutator with the Hamiltonian of mean force in eq. (62).
This is a promising approach, but the validity depends on the structure of $\varrho_{\beta}$ and the type of correlations between the system of interest and the environment. If there are no correlations $\varrho_{\beta}=\varrho_{S,\beta}\otimes\varrho_{E,\beta}$ this indeed holds since
@@
$=$ $-\frac{1}{\beta}[X_{S},H_{S}^{\ast}]$ (65)
We leave the question if this works out if classical and even quantum correlations are involved for future investigation. The interested reader can refer to Apx.C for an initial approach. Further, it remains to analyse whether the scalar product based on the transformation $\Sigma$ allows for a unique decomposition of the master equation (49). In ref. [37] it was established for the deformed scalar product eq. (54) with $\alpha=\frac{1}{2}$ that such a splitting exists but a proof for arbitrary $\alpha$ requires different techniques.
@@ VII Conclusion
We have suggested a definition of work performed on systems coupled to an environment, where the environment does not need to be in an equilibrium state and the coupling does not need to be weak. Our suggestion is based on projection operator techniques and the concept of minimal dissipation, which has recently been brought forward as a means to define work in open quantum systems [14, 34, 21]. For classical systems we show that the concept of minimal dissipation can be used to obtain a definition of work that has the correct limit for quasi-static processes performed on systems coupled to a heat bath. This is achieved by using an inner product with the global equilibrium distribution as weight. In contrast, the original proposal [14] established a splitting of the reduced dynamics based on an unweighted product and the minimization of the so called dissipative part. Our investigation suggests that the corresponding effective Hamiltonian is not directly related to thermodynamic work.
For systems and environments with equilibrium states that factorize, the ideas can directly be transferred from the classical to the quantum mechanical case. For more complex cases there does not seem to be a general procedure to define an appropriately weighted inner product and a projection operator. We assume that there is no unique definition, which yields the correct limit, and that the inner product has to be chosen case by case such that it is appropriate to a given system. The approach that we proposed admits a work definition in composite systems that equilibrate but is limited to situations where the Hamiltonian is time independent. As soon as the composite system is subject to external driving the drift term obtained with the Zwanzig projector does not admit a straightforward relation to a Hamiltonian of mean force [15, 41].
@@ Acknowledgements.
Data sharing does not apply to this article as no datasets were generated or analyzed during the current study.
@@ Appendix C Similarity Transformation and Kubo-relation
If this indeed is true, the next step would be the study of separable equilibrium states, i.e.  those that can be represented as a sum over product states $\varrho_{\beta}=\sum_{j}p_{j}\varrho_{S,j}\otimes\varrho_{E,j}$.

### v2

@@ VI Projection operator for quantum systems
It is not straightforward to transfer the reasoning for classical systems to quantum systems in a universally valid manner. In this section we will briefly discuss why this transfer is difficult in general, and then show systems and environments with equilibrium states that factorize are an exception.
For the classical Zwanzig projector eq. (28) we began the constructions with the introduction of the extended set of relevant variables $\{\psi_{\alpha}\}$. For a single relevant quantum observable $A$ the projectors onto its eigenstates can play the same role.
Suppose $A$ has a discrete spectrum
@@
Due to orthogonality, we have $\Pi_{j}\Pi_{k}=\delta_{jk}\Pi_{j}$, much the same as in eq. (27). Both $\mathcal{P}$ and $\mathcal{P}^{\dagger}$ are idempotent. The projected $\mathcal{P}^{\dagger}\rho$ is still a density matrix. It can be interpreted as the projection on the equilibrium state after a non-selective measurement of $A$. By construction, the expectation value of $A$ is preserved, i.e., $\text{Tr}(A\rho)=\text{Tr}(A\mathcal{P}^{\dagger}\rho)$.
@@
In addition, this construction does not produce an equation of motion for the reduced density matrix $\rho_{S}$ of a composite system. For a general density matrix, we find that $\rho_{S}=\text{Tr}_{E}(\rho)$ is not equal to $\text{Tr}_{E}(\mathcal{P}^{\dagger}\rho)$. Equality is only achieved if $\rho_{S}$ commutes with the relevant observable.
Instead, the projector can be built to project out all operators acting in $\mathcal{H}_{S}$. To implement this, we need a suitable scalar product similar to eq. (29). There is no unique quantum analogue, but an entire class of scalar products
@@
which give a weight to the commutativity property with $H$ [45]. A common choice in quantum statistical mechanics is to average over $\alpha$
@@
This scalar product is known under various names like Mori scalar product, Kubo’s canonical correlation, Bogoliubov inner product or Duhamel two-point function [33, 34, 46].
@@
The advantage of this transformation is, that
@@
where $\Sigma_{S}$ maps system operators $X_{S}\in\mathcal{H}_{S}$ to system operators: $\Sigma_{S}X_{S}=\text{Tr}_{E}(\Sigma X_{S})$. With these definitions, we formally have the desired $\mathcal{P}X_{S}=X_{S}$ and $\text{Tr}_{E}(\mathcal{P}^{\dagger}\rho)=\text{Tr}_{E}(\rho)=\rho_{S}$, i.e. all observables of the system are relevant and the relevant density yields the reduced system state if we trace over the environmental degrees of freedom.
@@
For the last step the relation eq. (61) has to hold when replacing $\Sigma\to\Sigma_{S}$ and $\varrho_{\beta}\to\varrho_{S,\beta}$, i.e.
$\frac{1}{\beta}\Sigma_{S}[X_{S},\ln\varrho_{S,\beta}]=\frac{1}{\beta}[X_{S},\varrho_{S,\beta}]$ (67)
has to be fulfilled. Under this condition the desired commutator with the Hamiltonian of mean force is obtained in eq. (66) from $\frac{1}{\beta}[X_{S},\rho_{S,\beta}]=-\Sigma_{S}[X_{S},H_{S}^{\ast}]$.
This is a promising approach, but the validity of eq. (67) depends on the structure of $\varrho_{\beta}$ and the type of correlations between the system of interest and the environment. If there are no correlations $\varrho_{\beta}=\varrho_{S,\beta}\otimes\varrho_{E,\beta}$ eq. (67) indeed holds since
@@
$=$ $-\frac{1}{\beta}[X_{S},\varrho_{S,\beta}]\;.$ (70)
We leave the question whether eq. (67) also holds if classical and even quantum correlations are involved for future investigation. The interested reader can refer to apx.C for an initial approach. Further, it remains to analyse whether the scalar product based on the transformation $\Sigma$ allows for a unique decomposition of the master equation (53). In ref. [44] it was established for the deformed scalar product eq. (58) with $\alpha=\frac{1}{2}$ that such a splitting exists but a proof for arbitrary $\alpha$ requires different techniques.
@@ VII Conclusion
We have suggested a definition of work performed on systems coupled to an environment, where the environment does not need to be in an equilibrium state and the coupling does not need to be weak. Our suggestion is based on projection operator techniques and the concept of minimal dissipation, which has recently been brought forward as a means to define work in open quantum systems [14, 41, 27]. For classical systems we show that the concept of minimal dissipation can be used to obtain a definition of work that has the correct limit for quasi-static processes performed on systems coupled to a heat bath. This is achieved by using an inner product with the global equilibrium distribution as weight. In contrast, the original proposal [14] established a splitting of the reduced dynamics based on an unweighted product and the minimization of the so-called dissipative part. Our investigation suggests that the corresponding effective Hamiltonian is not directly related to thermodynamic work.
For systems and environments with equilibrium states that factorize, the ideas can be transferred directly from the classical to the quantum mechanical case. For more complex cases there is no general procedure to define an appropriately weighted inner product and a projection operator. We assume that there is no unique definition. Thus, the inner product has to be chosen case by case such that it is appropriate to a given system. The approach that we proposed admits a work definition in composite systems that equilibrate but is limited to situations where the Hamiltonian is time independent. As soon as the composite system is subject to external driving the drift term obtained with the Zwanzig projector does not admit a straightforward relation to a Hamiltonian of mean force [25, 47].
@@ Acknowledgements.
Data sharing does not apply to this article as no datasets were generated or analysed during the current study.
@@ Appendix C Similarity Transformation and Kubo-relation
If this is indeed true, the next step would be the study of separable equilibrium states, i.e.  those that can be represented as a sum over product states $\varrho_{\beta}=\sum_{j}p_{j}\varrho_{S,j}\otimes\varrho_{E,j}$.